# LoRA: Low-Rank Adaptation of Large Language Models

### v1

@@ 7 Conclusion and Future Work
Fine-tuning enormous language models is prohibitively expensive in terms of both the hardware requirement and the storage/switching cost for hosting multiple instances. We propose LoRA, an efficient adaptation strategy that neither introduces inference latency nor reduces input sequence length while retaining model quality. Importantly, it allows for quick task-switching when deployed as a service by sharing the vast majority of the model parameters. While we focused on Transformer, the proposed principles are generally applicable to any neural networks with dense layers.
LoRA can potentially work in tandem with other fine-tuning techniques. In the future, we hope to explore only tuning some layers or adding adversarial training. Finally, the rank-deficiency of $\Delta W$ suggests that $W$ could be rank-deficient as well, which might inspire lots of future works.
Acknowledgments
We thank in alphabetical order Jianfeng Gao, Jade Huang, Jiayuan Huang, Lisa Xiang Li, Xiaodong Liu, Yabin Liu, Benjamin Van Durme, Luis Vargas, Haoran Wei, Peter Welinder, and Greg Yang for providing valuable feedback.
@@ Appendix A Large Language Models Still Need Parameter Updates
Few-shot learning, or prompt engineering, is very advantageous when we only have a handful of training samples. However, in practice, we can often afford to curate a few thousand or more training examples for performance-sensitive applications. As shown in Table 6, fine-tuning improves the model performance drastically compared to few-shot learning on datasets large and small. We take the GPT-3 few-shot result on RTE from the GPT-3 paper [6]. For MNLI-m task, we use two demonstrations per class or six in-context examples in total.
@@
Table 6: Fine-tuning significantly outperforms few-shot learning on GPT-3 [6].
Appendix B Dataset Details
MultiNLI is a natural language inference dataset  [37], consisting of $392,702$ training and $9,815$ validation examples. For pre-processing, we encode the context $x=\{[premise]:,\text{premise},[hypothesis]:,\text{hypothesis}\}$; target $y\in\{entailment,neutral,contradiction\}$. The dataset is mostly released under the OANC’s license, and subsets of it are under other permissive licenses such as the Creative Commons Share-Alike 3.0 Unported License.
@@ Appendix B Dataset Details
WikiSQL is introduced in  [40] and contains 56,35556,355/$8,421$ training/validation examples. The task is to generate SQL queries from natural language questions and table schemata. We encode context as $x=\{\text{table schema},\text{query}\}$ and target as $y=\{\text{SQL}\}$. The dataset is release under the BSD 3-Clause License.
@@
E2E NLG Challenge was first introduced in  [28] as a dataset for training end-to-end, data-driven natural language generation systems and is commonly used for data-to-text evaluation. The E2E dataset consists of 42K~42K training, 4.6K~4.6K validation, and 4.6K~4.6K test examples from the restaurant domain. Each source table used as input can have multiple references. Each sample input $(x,y)$ consists of a sequence of slot-value pairs, along with a corresponding natural language reference text. The dataset is released under Creative Commons BY-NC-SA 4.0.
@@
DART is an open-domain data-to-text dataset described in [27]. DART inputs are structured as sequences of ENTITY | RELATION | ENTITY triples. With 82K~82K examples in total, DART is a significantly larger and more complex data-to-text task compared to E2E. The dataset is released under the MIT license.
@@
WebNLG is another commonly used dataset for data-to-text evaluation [10]. With 22K~22K examples in total WebNLG comprises 14 distinct categories, nine of which are seen during training. Since five of the 14 total categories are not seen during training, but are represented in the test set, evaluation is typically broken out by “seen” categories (S), “unseen” categories (U) and “all” (A). Each input example is represented by a sequence of SUBJECT | PROPERTY | OBJECT triples. The dataset is released under Creative Commons BY-NC-SA 4.0.
Appendix C Hyperparameters Used in Experiments
C.1 GPT-3 Experiments
The training hyperparameters used in our GPT-3 experiments are listed in Table 7.
Hyperparameters Fine-Tune Prefix-Embed Prefix-Tune LoRA
@@ C.1 GPT-3 Experiments
Batch Size 128
# Epoch 2
Warmup Tokens 250,000
Learning Rate Schedule Linear
Learning Rate 5.00E-06 5.00E-04 1.00E-04 2.00E-04
Table 7: The training hyperparameters used for different GPT-3 adaption methods. We use the same hyperparameters for all datasets.
C.2 GPT-2 LoRA
The hyperparameters used for LoRA in GPT-2 are listed in Table 8. For those used for other baselines, see [21].
@@ C.2 GPT-2 LoRA
Appendix D Combining LoRA with Prefix Tuning
@@
LoRA+PrefixEmbed (LoRA+PE) combines LoRA with prefix-embedding tuning, where we insert $l_{p}+l_{i}$ special tokens whose embeddings are treated as trainable parameters. For more on prefix-embedding tuning, see Sec. 5.1.
LoRA+PrefixLayer (LoRA+PL) combines LoRA with prefix-layer tuning. We also insert $l_{p}+l_{i}$ special tokens; however, instead of letting the hidden representations of these tokens evolve naturally, we replace them after every Transformer block with an input agnostic vector. Thus, both the embeddings and subsequent Transformer block activations are treated as trainable parameters. For more on prefix-layer tuning, see Sec. 5.1.
@@
Appendix E Additional Task-Based Experiments
E.1 Additional Experiments on GPT-3
@@
Fine-Tune - 175B 73.0 89.5
PrefixEmbed $l_{p}=32,l_{i}=8$ 0.39 M 55.9 84.9
$l_{p}=64,l_{i}=8$ 0.88 M 58.7 88.1
$l_{p}=128,l_{i}=8$ 1.67 M 60.6 88.0
$l_{p}=256,l_{i}=8$ 3.24 M 63.1 88.6
$l_{p}=512,l_{i}=8$ 6.40 M 55.9 85.8
PrefixLayer $l_{p}=2,l_{i}=2$ 5.06 M 68.5 89.2
@@
E.2 Additional Experiments on GPT-2
We also repeat our experiment on DART [27] and WebNLG [10] following the setup of [21]. The result is shown in Table 10. Similar to our result on E2E NLG Challenge, reported in Sec. 5, LoRA performs better than or at least on-par with prefix-based approaches given the same number of trainable parameters.
Method # Trainable DART
[…]
GPT-2 Medium
Fine-Tune 354M 46.0($\pm 0.1$) 0.39 0.46
Adapter 10M 45.4($\pm 0.1$) 0.38 0.46
FT-Top2 24M 38.1($\pm 0.3$) 0.34 0.56
Prefix 0.35M 45.7($\pm 0.2$) 0.38 0.46
LoRA 0.35M 47.1($\pm 0.2$) 0.39 0.46
GPT-2 Large
Fine-Tune 774M 46.5($\pm 0.1$) 0.39 0.45
Prefix 0.77M 46.5($\pm 0.2$) 0.38 0.45
LoRA 0.77M 47.5($\pm 0.1$) 0.39 0.45
Table 10: GPT-2 with different adaptation methods on DART. The variances of MET and TER are around $0.01$ for all different adaption approaches.
Method WebNLG
[…]
U S A U S A U S A
GPT-2 Medium
Fine-Tune (354M) 30.4($\pm.5$) 63.2($\pm.3$) 47.6($\pm.4$) .32 .45 .39 .69 .34 .50
Adapter (10M) 47.9($\pm.2$) 61.1($\pm.4$) 55.2($\pm.3$) .38 .43 .41 .45 .35 .39
FT-Top2 (24M) 13.7($\pm.6$) 50.1($\pm.4$) 33.5 ($\pm.4$) .16 .35 .26 1.03 .52 .75
Prefix (0.35M) 44.1($\pm.2$) 63.1($\pm.1$) 54.4($\pm.1$) .37 .45 .41 .50 .34 .41
LoRA (0.35M) 46.7($\pm.4$) 62.1($\pm.2$) 55.3($\pm.2$) .38 .44 .41 .46 .33 .39
GPT-2 Large
Fine-Tune (774M) 41.7($\pm.5$) 64.6($\pm.4$) 54.2($\pm.4$) .37 .46 .42 .56 .33 .43
Prefix (0.77M) 47.0($\pm.2$) 64.2($\pm.4$) 56.4($\pm.1$) .39 .45 .42 .49 .33 .40
LoRA (0.77M) 48.4($\pm.3$) 64.0($\pm.3$) 57.0($\pm.1$) .39 .45 .42 .45 .32 .38
Table 11: GPT-2 with different adaptation methods on WebNLG. The variances of MET and TER are around $0.01$ variance for all different adaption approaches. “U” indicates unseen categories, “S” indicates seen categories, and “A” indicates all categories in the test set of webNLG.
E.3 Low-Data Regime
@@
Appendix F Measuring Similarity Between Subspaces
In this paper we use the measure $\phi(A,B,i,j)=\psi(U_{A}^{i},U_{B}^{j})=\frac{\|U_{A}^{i\top}U_{B}\|_{F}^{2}}{\min\{i,j\}}$ to measure the subspace similarity between two column orthonormal matrices $U_{A}^{i}\in\mathbb{R}^{d\times i}$ and $U_{B}^{j}\in\mathbb{R}^{d\times j}$, obtained by taking columns of the left singular matrices of $A$ and $B$. We point out that this similarity is simply a reverse of the standard Projection Metric that measures distance between subspaces [dist].
To be concrete, let the singular values of $U_{A}^{i\top}U_{B}^{j}$ to be $\sigma_{1},\sigma_{2},\cdots,\sigma_{p}$ where $p=\min\{i,j\}$. We know that the Projection Metric [dist] is defined as:

### v2

@@ 8 Conclusion and Future Work
Fine-tuning enormous language models is prohibitively expensive in terms of the hardware required and the storage/switching cost for hosting independent instances for different tasks. We propose LoRA, an efficient adaptation strategy that neither introduces inference latency nor reduces input sequence length while retaining high model quality. Importantly, it allows for quick task-switching when deployed as a service by sharing the vast majority of the model parameters. While we focused on Transformer language models, the proposed principles are generally applicable to any neural networks with dense layers.
There are many directions for future works. 1) LoRA can be combined with other efficient adaptation methods, potentially providing orthogonal improvement. 2) The mechanism behind fine-tuning or LoRA is far from clear – how are features learned during pre-training transformed to do well on downstream tasks? We believe that LoRA makes it more tractable to answer this than full fine-tuning. 3) We mostly depend on heuristics to select the weight matrices to apply LoRA to. Are there more principled ways to do it? 4) Finally, the rank-deficiency of $\Delta W$ suggests that $W$ could be rank-deficient as well, which can also be a source of inspiration for future works.
@@ Appendix A Large Language Models Still Need Parameter Updates
Few-shot learning, or prompt engineering, is very advantageous when we only have a handful of training samples. However, in practice, we can often afford to curate a few thousand or more training examples for performance-sensitive applications. As shown in Table 8, fine-tuning improves the model performance drastically compared to few-shot learning on datasets large and small. We take the GPT-3 few-shot result on RTE from the GPT-3 paper (Brown et al., 2020). For MNLI-matched, we use two demonstrations per class and six in-context examples in total.
@@
Table 8: Fine-tuning significantly outperforms few-shot learning on GPT-3 (Brown et al., 2020).
Appendix B Inference Latency Introduced by Adapter Layers
Adapter layers are external modules added to a pre-trained model in a sequential manner, whereas our proposal, LoRA, can be seen as external modules added in a parallel manner. Consequently, adapter layers must be computed in addition to the base model, inevitably introducing additional latency. While as pointed out in Rücklé et al. (2020), the latency introduced by adapter layers can be mitigated when the model batch size and/or sequence length is large enough to full utilize the hardware parallelism. We confirm their observation with a similar latency study on GPT-2 medium and point out that there are scenarios, notably online inference where the batch size is small, where the added latency can be significant.
We measure the latency of a single forward pass on an NVIDIA Quadro RTX8000 by averaging over 100 trials. We vary the input batch size, sequence length, and the adapter bottleneck dimension $r$. We test two adapter designs: the original one by Houlsby et al. (2019), which we call $\text{Adapter}^{\text{H}}$, and a recent, more efficient variant by Lin et al. (2020), which we call $\text{Adapter}^{\text{L}}$. See Section 5.1 for more details on the designs. We plot the slow-down in percentage compared to the no-adapter baseline in Figure 5.
Figure 5: Percentage slow-down of inference latency compared to the no-adapter ($r=0$) baseline. The top row shows the result for $\text{Adapter}^{\text{H}}$ and the bottom row $\text{Adapter}^{\text{L}}$. Larger batch size and sequence length help to mitigate the latency, but the slow-down can be as high as over 30% in an online, short-sequence-length scenario. We tweak the colormap for better visibility.
Appendix C Dataset Details
GLUE Benchmark is a wide-ranging collection of natural language understanding tasks. It includes MNLI (inference, Williams et al. (2018)), SST-2 (sentiment analysis, Socher et al. (2013)), MRPC (paraphrase detection, Dolan & Brockett (2005)), CoLA (linguistic acceptability, Warstadt et al. (2018)), QNLI (inference, Rajpurkar et al. (2018)), QQP [^8] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs (question-answering), RTE (inference), and STS-B (textual similarity, Cer et al. (2017)). The broad coverage makes GLUE benchmark a standard metric to evaluate NLU models such as RoBERTa and DeBERTa. The individual datasets are released under different permissive licenses.
@@ Appendix C Dataset Details
WikiSQL is introduced in  Zhong et al. (2017) and contains 56,35556,355/$8,421$ training/validation examples. The task is to generate SQL queries from natural language questions and table schemata. We encode context as $x=\{\text{table schema},\text{query}\}$ and target as $y=\{\text{SQL}\}$. The dataset is release under the BSD 3-Clause License.
[…]
@@
E2E NLG Challenge was first introduced in  Novikova et al. (2017) as a dataset for training end-to-end, data-driven natural language generation systems and is commonly used for data-to-text evaluation. The E2E dataset consists of roughly $42,000$ training, $4,600$ validation, and $4,600$ test examples from the restaurant domain. Each source table used as input can have multiple references. Each sample input $(x,y)$ consists of a sequence of slot-value pairs, along with a corresponding natural language reference text. The dataset is released under Creative Commons BY-NC-SA 4.0.
@@
DART is an open-domain data-to-text dataset described in Nan et al. (2020). DART inputs are structured as sequences of ENTITY — RELATION — ENTITY triples. With 82K~82K examples in total, DART is a significantly larger and more complex data-to-text task compared to E2E. The dataset is released under the MIT license.
@@
WebNLG is another commonly used dataset for data-to-text evaluation (Gardent et al., 2017). With 22K~22K examples in total WebNLG comprises 14 distinct categories, nine of which are seen during training. Since five of the 14 total categories are not seen during training, but are represented in the test set, evaluation is typically broken out by “seen” categories (S), “unseen” categories (U) and “all” (A). Each input example is represented by a sequence of SUBJECT — PROPERTY — OBJECT triples. The dataset is released under Creative Commons BY-NC-SA 4.0.
Appendix D Hyperparameters Used in Experiments
D.1 RoBERTa
We train using AdamW with a linear learning rate decay schedule. We sweep learning rate, number of training epochs, and batch size for LoRA. Following Liu et al. (2019), we initialize the LoRA modules to our best MNLI checkpoint when adapting to MRPC, RTE, and STS-B, instead of the usual initialization; the pre-trained model stays frozen for all tasks. We report the median over 5 random seeds; the result for each run is taken from the best epoch. For a fair comparison with the setup in Houlsby et al. (2019) and Pfeiffer et al. (2021), we restrict the model sequence length to 128 and used a fixed batch size for all tasks. Importantly, we start with the pre-trained RoBERTa large model when adapting to MRPC, RTE, and STS-B, instead of a model already adapted to MNLI. The runs with this restricted setup are marked with $\dagger$. See the hyperparameters used in our runs in Table 9.
Method Dataset MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
@@ D.1 RoBERTa
Warmup Ratio 0.06
LR Schedule Linear
RoBERTa base LoRA Batch Size 16 16 16 32 32 16 32 16
# Epochs 30 60 30 80 25 25 80 40
Learning Rate 5E-04 5E-04 4E-04 4E-04 4E-04 5E-04 5E-04 4E-04
LoRA Config. $r_{q}=r_{v}=8$
LoRA $\alpha$ 8
Max Seq. Len. 512
RoBERTa large LoRA Batch Size 4 4 4 4 4 4 8 8
# Epochs 10 10 20 20 10 20 20 30
Learning Rate 3E-04 4E-04 3E-04 2E-04 2E-04 3E-04 4E-04 2E-04
LoRA Config. $r_{q}=r_{v}=8$
LoRA $\alpha$ 16
Max Seq. Len. 128 128 512 128 512 512 512 512
[…]
# Epochs 10 10 20 20 10 20 20 10
Learning Rate 3E-04 4E-04 3E-04 2E-04 2E-04 3E-04 4E-04 2E-04
LoRA Config. $r_{q}=r_{v}=8$
LoRA $\alpha$ 16
Max Seq. Len. 128
[…]
# Epochs 10 20 20 20 10 20 20 20
Learning Rate 3E-05 3E-05 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04
Bottleneck $r$ 64
Max Seq. Len. 128
[…]
# Epochs 5 20 20 20 10 20 20 20
Learning Rate 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04
Bottleneck $r$ 16
Max Seq. Len. 128
[…]
# Epochs 10 5 10 10 5 20 20 10
Learning Rate 3E-05 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04
Bottleneck $r$ 64
Max Seq. Len. 128
[…]
# Epochs 10 5 10 10 5 20 20 10
Learning Rate 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04 3E-04
Bottleneck $r$ 8
Max Seq. Len. 128
Table 9: The hyperparameters we used for RoBERTa on the GLUE benchmark.
D.2 DeBERTa
We again train using AdamW with a linear learning rate decay schedule. Following He et al. (2021), we tune learning rate, dropout probability, warm-up steps, and batch size. We use the same model sequence length used by (He et al., 2021) to keep our comparison fair. Following He et al. (2021), we initialize the LoRA modules to our best MNLI checkpoint when adapting to MRPC, RTE, and STS-B, instead of the usual initialization; the pre-trained model stays frozen for all tasks. We report the median over 5 random seeds; the result for each run is taken from the best epoch. See the hyperparameters used in our runs in Table 10.
Method Dataset MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Optimizer AdamW
Warmup Ratio 0.1
LR Schedule Linear
DeBERTa XXL LoRA Batch Size 8 8 32 4 6 8 4 4
# Epochs 5 16 30 10 8 11 11 10
Learning Rate 1E-04 6E-05 2E-04 1E-04 1E-04 1E-04 2E-04 2E-04
Weight Decay 0 0.01 0.01 0 0.01 0.01 0.01 0.1
CLS Dropout 0.15 0 0 0.1 0.1 0.2 0.2 0.2
LoRA Config. $r_{q}=r_{v}=8$
LoRA $\alpha$ 8
Max Seq. Len. 256 128 128 64 512 320 320 128
Table 10: The hyperparameters for DeBERTa XXL on tasks included in the GLUE benchmark.
D.3 GPT-2
We train all of our GPT-2 models using AdamW (Loshchilov & Hutter, 2017) with a linear learning rate schedule for 5 epochs. We use the batch size, learning rate, and beam search beam size described in Li & Liang (2021). Accordingly, we also tune the above hyperparameters for LoRA. We report the mean over 3 random seeds; the result for each run is taken from the best epoch. The hyperparameters used for LoRA in GPT-2 are listed in Table 11. For those used for other baselines, see Li & Liang (2021).
@@ D.3 GPT-2
D.4 GPT-3
For all GPT-3 experiments, we train using AdamW (Loshchilov & Hutter, 2017) for 2 epochs with a batch size of 128 samples and a weight decay factor of 0.1. We use a sequence length of 384 for WikiSQL (Zhong et al., 2017), 768 for MNLI (Williams et al., 2018), and 2048 for SAMSum (Gliwa et al., 2019). We tune learning rate for all method-dataset combinations. See Section D.4 for more details on the hyperparameters used. For prefix-embedding tuning, we find the optimal $l_{p}$ and $l_{i}$ to be 256 and 8, respectively, totalling $3.2M$ trainable parameters. We use $l_{p}=8$ and $l_{i}=8$ for prefix-layer tuning with $20.2M$ trainable parameters to obtain the overall best performance. We present two parameter budgets for LoRA: 4.7M ($r_{q}=r_{v}=1$ or $r_{v}=2$) and 37.7M ($r_{q}=r_{v}=8$ or $r_{q}=r_{k}=r_{v}=r_{o}=2$). We report the best validation performance from each run. The training hyperparameters used in our GPT-3 experiments are listed in Table 12.
Hyperparameters Fine-Tune PreEmbed PreLayer BitFit $\text{Adapter}^{\text{H}}$ LoRA
Optimizer AdamW
Batch Size 128
# Epoch 2
Warmup Tokens 250,000
LR Schedule Linear
Learning Rate 5.00E-06 5.00E-04 1.00E-04 1.6E-03 1.00E-04 2.00E-04
Table 12: The training hyperparameters used for different GPT-3 adaption methods. We use the same hyperparameters for all datasets after tuning learning rate.
Appendix E Combining LoRA with Prefix Tuning
@@
LoRA+PrefixEmbed (LoRA+PE) combines LoRA with prefix-embedding tuning, where we insert $l_{p}+l_{i}$ special tokens whose embeddings are treated as trainable parameters. For more on prefix-embedding tuning, see Section 5.1.
LoRA+PrefixLayer (LoRA+PL) combines LoRA with prefix-layer tuning. We also insert $l_{p}+l_{i}$ special tokens; however, instead of letting the hidden representations of these tokens evolve naturally, we replace them after every Transformer block with an input agnostic vector. Thus, both the embeddings and subsequent Transformer block activations are treated as trainable parameters. For more on prefix-layer tuning, see Section 5.1.
@@
Appendix F Additional Empirical Experiments
F.1 Additional Experiments on GPT-2
We also repeat our experiment on DART (Nan et al., 2020) and WebNLG (Gardent et al., 2017) following the setup of Li & Liang (2021). The result is shown in Table 13. Similar to our result on E2E NLG Challenge, reported in Section 5, LoRA performs better than or at least on-par with prefix-based approaches given the same number of trainable parameters.
Method # Trainable DART
[…]
GPT-2 Medium
Fine-Tune 354M 46.2 0.39 0.46
$\text{Adapter}^{\text{L}}$ 0.37M 42.4 0.36 0.48
$\text{Adapter}^{\text{L}}$ 11M 45.2 0.38 0.46
$\text{FT}^{\text{Top2}}$ 24M 41.0 0.34 0.56
PrefLayer 0.35M 46.4 0.38 0.46
[…]
GPT-2 Large
Fine-Tune 774M 47.0 0.39 0.46
[…]
PrefLayer 0.77M 46.7 0.38 0.45
[…]
Table 13: GPT-2 with different adaptation methods on DART. The variances of MET and TER are less than $0.01$ for all adaption approaches.
Method WebNLG
[…]
U S A U S A U S A
GPT-2 Medium
Fine-Tune (354M) 27.7 64.2 46.5 .30 .45 .38 .76 .33 .53
$\text{Adapter}^{\text{L}}$ (0.37M) 45.1 54.5 50.2 .36 .39 .38 .46 .40 .43
$\text{Adapter}^{\text{L}}$ (11M) 48.3 60.4 54.9 .38 .43 .41 .45 .35 .39
$\text{FT}^{\text{Top2}}$ (24M) 18.9 53.6 36.0 .23 .38 .31 .99 .49 .72
Prefix (0.35M) 45.6 62.9 55.1 .38 .44 .41 .49 .35 .40
[…]
GPT-2 Large
Fine-Tune (774M) 43.1 65.3 55.5 .38 .46 .42 .53 .33 .42
[…]
Prefix (0.77M) 47.7 63.4 56.3 .39 .45 .42 .48 .34 .40
[…]
Table 14: GPT-2 with different adaptation methods on WebNLG. The variances of MET and TER are less than $0.01$ for all the experiments we ran. “U” indicates unseen categories, “S” indicates seen categories, and “A” indicates all categories in the test set of WebNLG.
F.2 Additional Experiments on GPT-3
@@
Fine-Tune - 175B 73.8 89.5
PrefixEmbed $l_{p}=32,l_{i}=8$ 0.4 M 55.9 84.9
$l_{p}=64,l_{i}=8$ 0.9 M 58.7 88.1
$l_{p}=128,l_{i}=8$ 1.7 M 60.6 88.0
$l_{p}=256,l_{i}=8$ 3.2 M 63.1 88.6
$l_{p}=512,l_{i}=8$ 6.4 M 55.9 85.8
PrefixLayer $l_{p}=2,l_{i}=2$ 5.1 M 68.5 89.2
@@
F.3 Low-Data Regime
@@
Appendix G Measuring Similarity Between Subspaces
In this paper we use the measure $\phi(A,B,i,j)=\psi(U_{A}^{i},U_{B}^{j})=\frac{\|U_{A}^{i\top}U_{B}\|_{F}^{2}}{\min\{i,j\}}$ to measure the subspace similarity between two column orthonormal matrices $U_{A}^{i}\in\mathbb{R}^{d\times i}$ and $U_{B}^{j}\in\mathbb{R}^{d\times j}$, obtained by taking columns of the left singular matrices of $A$ and $B$. We point out that this similarity is simply a reverse of the standard Projection Metric that measures distance between subspaces Ham & Lee (2008).
To be concrete, let the singular values of $U_{A}^{i\top}U_{B}^{j}$ to be $\sigma_{1},\sigma_{2},\cdots,\sigma_{p}$ where $p=\min\{i,j\}$. We know that the Projection Metric Ham & Lee (2008) is defined as: